What are the missing terms of the geometric sequence -3, _, _, -375?
A)-127, -251
B) 127, 251
C)15, 75
D) -15, -75

Answer by DrBeeee(684)

A geometric series means that each subsequent term is the product of the term and a CONSTANT.
Let n be the constant.
The given sequence is
(1) -3, _, _, -375 or
(2) -3, -3*n, -3*n*n, -3*n*n*n
Note that the fourth term is
(3) -3*n*n*n = -375 or
(4) 3*n^3 = 375 or
(5) n^3 = 125 or
(6) n = 5
The complete sequence is
(7) -3, -15, -75, -375
Answer: The missing terms are -15 and -75, choice D.